# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

### v1

@@ 5 Resource and efficiency
Table  6 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. Due to the structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage by 5GB compared to LoRA. Additionally, Bone’s token throughput is only two-thirds of LoRA’s. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput.
Despite the fact that the Bone structure significantly outperforms LoRA variants in fine-tuning, we cannot ignore the efficiency issues. Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
@@ 6 Conclusion
This paper introduces a new PEFT technique called Bone(Block Affine), which employs block affine transformations on the weight matrix to capture the internal correlations within $W$, while also avoiding the complexity of initialization. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.

### v2

@@ 5 Resource and efficiency
Table  7 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. The structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput. At the end of the table, we provide the real-world training scenarios comparing the memory usage and token throughput of Bone and LoRA.
Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
@@
RWKV-3B LoRA(model checkpoint) 55.1M 4*15328 MB 15.6 kt/s
Bone(model checkpoint) 55.1M 4*15305 MB 14.2 kt/s
@@ 6 Conclusion
This paper introduces a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information. Based on this theory, we designed a new PEFT technique called Bone (Block Affine). Additionally, we employed block matrix multiplication to efficiently promote information fusion within the weights in the Bone structure. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.